# Spin Response of a Magnetic Monopole and Quantum Hall Response in Topological Lattice Models through Local Invariants and Light

### v1

@@ Abstract
Here, we elaborate on and develop the geometrical approach introduced in K. Le Hur, Physics Reports 1104 1-42 (2025) between the magnetic monopole created from a radial field, quantum physics and topological lattice models through quantum phase transitions. We introduce an effective magnetic moment for a monopole when applying an additional source field along z-direction which also mediates the quantum phase transition. We present its relation with the transverse pumped quantum Hall current. The magnetic susceptibility can be introduced as a measure of the topological invariant i.e. remains quantized within the topological phase until the transition. We show the relation with two-dimensional topological lattice models such as a honeycomb Haldane model in real space. We develop the theory and present a numerical analysis between local invariants in momentum space introduced from Dirac points, correlation functions and the responses to circularly polarized light. We develop the formalism for coupled-planes materials including the possibility of quantum spin Hall effect and address a relation between the Ramanujan infinite alternating series and an interface in real space with a topological number one-half.
@@ 1 Introduction
The theory of magnetic monopoles attracts attention since the works of P. Curie [1] and P. A. M. Dirac [2, 3]. Detecting such monopoles associated to the planetary system, the universe, is yet of great interest. Recently, one of us has introduced a geometrical approach showing a correspondence between classical magnetic monopoles and quantum magnetic monopoles from a formulation at the poles on the sphere [4]. The magnetic monopole is produced from a radial magnetic field. Topological properties or equivalently the magnetic topological charge are generally measured through the integral of the magnetic field or through the integral of the Berry curvature in quantum physics on the surface area [5]. This approach [4] can be precisely rephrased into an effective cartesian metric in the vicinity of the poles in terms of a classical vector potential or in terms of the Berry gauge potential [5] in quantum physics associated to a coherent gauge (phase) representation of the eigenstates for the spin-1/2 particle on the Bloch sphere. Equivalently, the information of the magnetic charge is transported on each side of the equator to the poles in a thin handle. This formalism finds several practical applications [4]. It allows for a description of the quantum Hall response when building an analogy to the Newtonian approach. Applying an electric field along the polar angle an electron goes down from north to south poles with the induction of a transverse pumped current or a topological transverse pumped charge which is related to the quantum Hall conductivity and to the physics at the edges on a cylinder. It also allows us to show that photo-electric effects can become topological [6]. The responses to circularly polarized light can measure the topological charge or the quantum Hall conductivity from the poles which can be directly applied in two-dimensional (2D) topological lattice models through the Haldane model [7] and the quantum anomalous Hall effect (QAH) [8, 9] associated to the $\mathbb{Z}$ quantized invariant, where the poles correspond to the two Dirac points of the honeycomb lattice. The quantum Hall conductivity [10] can be equivalently measured through a summation on the wavevectors of the whole Brillouin zone from the same light signals, as introduced in Refs. [11, 12, 13]. Our geometrical formulation can also account for interaction effects in the detection of light in a quantitative manner through a variational stochastic approach [14, 15]. When adding interaction effects on the lattice, the variational stochastic approach is very efficient to reveal the first-order nature of the quantum phase transition towards the Mott phase [14]. This approach is also adapted for two-dimensional topological insulators, i.e. the quantum spin Hall (QSH) effect and the Kane-Mele model [16, 17, 18], where circularly polarized light can also detect the $\mathbb{Z}_{2}$ invariant locally from the Dirac points[6]. It is important to emphasize that magnetic monopoles through the effect of a radial magnetic field are realized in quantum circuits with applications e.g. to the Haldane model [19, 20, 21]. The quantum metric tensor is also measured associated to phase transitions [21]. Within our approach, we have shown that circularly polarized light produces the same effect as a boost of the azimuthal angle building a correspondence between quantum metric and responses to circularly polarized light through the introduction of a geometrical function [4]. It is also possible to measure the geometrical information through the optical conductivity in quantum materials [22].
@@
In this article, we develop our geometrical approach towards quantum phase transitions, addressing transport properties and responses to circularly polarized light. The quantum phase transition is induced through a fixed magnetic field $M$ polarized along $z$ direction. In previous works, we have shown how the topological invariant can also be re-written in terms of the spin responses at the poles related to a protected quantum dynamo effect [23, 4, 24], with applications on topological quantum wires [25, 26, 27]. In this work, we evaluate the spin response on the surface of the ball, which is similar to a Skyrmion for $M=0$ as shown in Fig. 1. We introduce the effective magnetic moment of the monopole as a function of $M$ associated to the spin response. We show how this is related to the transport properties. We introduce the associated susceptibility for the monopole which remains invariant within the topological phase until the phase transition and therefore can be introduced as a topological marker. To the best of our knowledge, this magnetic susceptibility of the monopole was not mentioned before in the literature. The magnetic moment has an orbital interpretation in the sense of the angular momentum associated to the transverse pumped current (charge) and this topological susceptibility shows then an analogy with orbital magnetism where the quantum Hall response is also thought of as a topological orbital susceptibility [28] corresponding to the derivative of the orbital magnetic moment with respect to the applied perpendicular magnetic field. We will then develop the theory for the Haldane model in two dimensions [7] with the introduction of a Semenoff mass [29] which corresponds e.g. to the effect of a charge density wave substrate [30]. Topological properties describing the occurrence of an edge mode are generally robust until the phase transition. Above the transition, particles progressively reside on one site preferably as a result of the proximity effect referring to a trivial insulator which does not present an edge mode and is characterized through a zero index (invariant). Finding observables to describe such a transition towards this trivial phase is yet of great interest e.g. related to the relative populations on the two sublattices. We take time to develop a numerical analysis showing how our geometrical approach from the poles and from the Dirac points is particularly efficient. The Berry curvatures at the two Dirac points encode information on the quantum Hall conductivity locally in momentum space [6]. We develop the analysis in real space related to the magnetic moment of the monopole and to correlation functions. It is relevant to mention here recent progress on measuring locally topological or geometrical properties in momentum space from photo-luminescence related to the Bloch sphere approach [31].
@@
The organization of our thoughts in this article is as follows. In Sec. 2, we develop the geometrical aspects of the monopole in quantum physics associated to the quantum phase transition and we present the evaluation of the spin response and of the effective magnetic moment as a function of the magnetic field $M$. We also present the relation with the (mean) quantized transverse Hall current from a Newtonian approach. In Sec. 3, we develop the theory for the honeycomb Haldane model in the presence of a Semenoff mass producing the quantum phase transition. We present a precise map from the rhomboid Brillouin zone onto the sphere and derive analytically and numerically the local responses associated to local invariants in momentum (reciprocal) space. The spin-1/2 on the Bloch sphere will correspond to the pseudo spin-1/2 measuring the population imbalance between the two sublattices of the honeycomb lattice. We introduce a real space analysis related to the effective magnetic moment of the monopole and address correlation functions numerically. We elaborate on the local responses from circularly polarized light associated to the quantum phase transition. In Sec. 4, we develop the theory of coupled-planes related to the $\frac{1}{2}$ topological invariant building an analogy between a topological interface and the Ramanujan infinite alternating series. We analyse transport properties and the responses to circularly light within this correspondence. The coupled-planes model includes the alternating presence of QAH and QSH effects. In Sec. 5, we summarize our findings. Appendices are devoted to additional mathematical proofs. In Appendix A, we elaborate on the relation between the geometrical approach and transport, quantum Hall response with a map onto a cylinder. In Appendix B, we present a geometrical justification of the $\frac{1}{2}$ topological invariant for the coupled planes when reaching the infinity limit.
@@ 2 Bloch Sphere as a Magnetic Monopole, Quantum Hall Response and Topological Magnetism
In this Section, we study the physical responses as a function of the parameter $M$ corresponding to the fixed additional magnetic field along $z$ direction which will drive the quantum phase transition associated to the Poincaré-Hopf theorem. We evaluate the local spin magnetization responses at and in the vicinity of the poles on the Bloch sphere as a function of $M$ related to topological properties produced by the radial magnetic field. We present a simple proof in a dressed polar angle representation showing how the topological number is generally measurable locally through the spin response [23, 4, 24]. We will then introduce the mean (medium) spin magnetization on the surface area and address the relation with the transverse quantum Hall current or transverse pumped charge in a Newtonian (gravitational) correspondence when driving a charge from north to south pole through a longitudinal electric field. This will give rise to the effective topological magnetic moment $\kappa(M)$, associated to the loops of transverse currents formed along the equatorial plane on the whole surface. When $M=0$, as a result of the radial magnetic field, magnetic moments of the two hemispheres compensate each other and $\kappa(M=0)=0$. For 2D topological lattice models, this will be compared with the local pseudo-spin response in real space. On the honeycomb lattice, we emphasize that the pseudo-spin will refer to the difference of occupancy between $A$ and $B$ sublattices. This analysis will also be related to the cylinder geometry [4].
@@
The last equation is equivalent to $|\alpha|^{2}=\cos^{2}\frac{\tilde{\theta}}{2}$ and $|\beta|^{2}=\sin^{2}\frac{\tilde{\theta}}{2}$. The topological invariant is measured for any fixed value of $M$. In this way, we verify the applicability of the local formulation of the topological invariant from the poles, which is gauge invariant and can be shown from general geometrical thoughts [4, 24] (see Appendix A). All the dependence on the parameter $M$ is then hidden into the definitions of the angles $\tilde{\theta}(\theta=0)$ and $\tilde{\theta}(\theta=\pi)$. We emphasize that the eigenstates are introduced within the same coherent gauge on the whole surface.
Figure 2: (Left) Classical magnetic monopole. We generalize the classical analogue of $A_{\varphi}$ introduced in Eq. (10) of Ref. [4] when including the magnetic field $M$ along $z$ direction as $A_{\varphi}=-\frac{\cos\theta}{2}-\frac{M}{4}\cos(2\theta)$ (for a unit sphere). We verify in this way that the topological charge remains protected for the classical monopole due to the presence of $M$ as it produces a component to the Berry gauge potential such that $A_{\varphi}^{M}(0)=A_{\varphi}^{M}(\pi)$. (Right) Representation of the effective topological properties for the quantum situation when including the magnetic field $M$ through the (dressed) angle $\tilde{\theta}(\theta=\pi)$ as a function of $M$. For all the phases the north pole is identified with $\theta=0$ i.e. $\tilde{\theta}=0$. Within the topological phase $B<M$, the south pole corresponds to $\tilde{\theta}(\theta=\pi)$ and the magnetic field is applied on the whole (dressed) sphere. For $M>B$, the north and south poles are identical and the equivalent geometry does not encircle the topological charge at the origin. For $M=B$, geometrical properties are equivalent to a half sphere corresponding to a half monopole or a half Skyrmion.
As shown in Fig. 2, for all values of $M$, the north pole associated to $\theta=0$ is also equivalent to $\tilde{\theta}=0$ i.e. $\tilde{\theta}(\theta=0)=0$. On the other hand, we do observe a quantum phase transition when $M=B$ from the angle $\tilde{\theta}$ : as long as $M<B$, then $\tilde{\theta}=\pi$ also corresponds to south pole i.e. to $\theta=\pi$ such that $\tilde{\theta}(\theta=\pi)=\pi$, when $M=B$ the topological properties become equivalent to half of a sphere $\tilde{\theta}(\theta=\pi)=\frac{\pi}{2}$ and for $M>B$ we also have the important equality $\tilde{\theta}(\theta=\pi)=\tilde{\theta}(\theta=0)$ implying $C=0$. The topological number then jumps from one when $M<B$ to zero when $M>B$. Within this formulation, at the quantum phase transition we also identify
@@
which can be interpreted as a half Skyrmion. In the topological phase, the physics is similar to a Skyrmion ; this correspondence will be re-inforced with the magnetic response or spin structure on the surface area in Fig. 1. It is interesting to mention that $A_{\varphi}(\tilde{\theta}(\theta=0))=+\frac{1}{2}$ corresponds to a topological characterization of a $\pi$ Berry phase encircling the north pole [24, 33]. The picture of the topological transition also agrees with the Poincaré-Hopf theorem and the halved surface area participating in the topological (geometrical) response in the dressed angle representation is similar to say that at the transition the topological charge also leaks out in Fig. 2.
We also have a correspondence with the spin response such that from Ehrenfest’s theorem $\langle\psi|\sigma_{z}|\psi\rangle=\langle\sigma_{z}\rangle=\cos\tilde{\theta}$. Then, the topological number can be reformulated through the spin polarizations at the north pole and at the dressed angle associated to $\theta=\pi$ :
@@
This definition agrees with previous results on one sphere when $M=0$ [23]. We emphasize that this formula was also adapted for the analysis of topological quantum wires with applications in real space associated to correlation functions [25, 26, 27]. The spin response is shown in Fig. 1 associated to the angle $\theta\in[0;\pi]$.
The profile of the magnetic structure at the poles then clearly reveals the quantum phase transition. At the topological quantum phase transition, we also have the identity
@@
For two spheres, it is possible to reach fractional $\frac{1}{2}$ numbers on a line or in a phase of the parameter space associated to the interaction between spins [4]. For $M>B$, the spin magnetizations associated to $\theta=0$ and $\theta=\pi$ are the same such that $C=0$ i.e. $\tilde{\theta}(\theta=0)=\tilde{\theta}(\theta=\pi)$.
@@
The pumped charge at an angle $\tilde{\theta}$ reads
@@
In this formulation, in the presence of the radial magnetic field acting on the spin of an electron and in the presence of the magnetic field source $M$, we measure the transverse pumped charge associated to an electric field applied along the polar angle such that $\hbar\dot{\tilde{\theta}}=eE$. At time $t=0$, a charge is present at north pole. The charge goes down similar to the apple and the gravitational force is a result of the electric field directed along the polar angle direction i.e. corresponds to a Coulomb force. The electron moves on the surface of this metallic ball associated to the Bloch sphere associated to the spin sector. The final time $T$ is fixed to be identical for any $M$ such that the angle $\tilde{\theta}$ reaches $\pi$ i.e. $T=\frac{eE}{\hbar\pi}=\frac{2eE}{h}$. Within the topological phase, the transverse pumped charge at south pole precisely measures the topological invariant itself $C=1$. When we reach the transition maintaining the same value of the (final) time $T$ for the measure, associated to Fig. 2, this corresponds to maintain the Berry gauge potential fixed to $A_{\varphi}(\tilde{\theta}=\pi)=A_{\varphi}(\tilde{\theta}=\frac{\pi}{2})=0$ for dressed angles $\tilde{\theta}\in[\frac{\pi}{2},\pi]$ such that there is no Berry curvature effectively present in the south hemisphere. In that case, the topological invariant becomes halved i.e. equal to $C_{1/2}$ and the transverse pumped charge is also halved. Above the transition, the charge remains at the north pole i.e. $A_{\varphi}(\tilde{\theta}=\pi)=A_{\varphi}(\tilde{\theta}=0)$ and there is no transverse pumped current. In Appendix A, we also show the relation with the cylinder geometry.
@@
Here, $\kappa$ is the mean of the spin response or the effective spin response as a function of $M$ associated to the surface area in Fig. 1. When $M=0$, from symmetry between the two hemispheres $\kappa=0$. In response to the radial magnetic field in Fig. 1 the two hemispheres produce an effective moment equal in amplitude but opposite in directions. We are precisely questioning the (topological) properties of $\kappa$ as a function of $M$ inserting the equation
@@
When $M\rightarrow+\infty$ this results in $\langle\sigma_{z}\rangle\rightarrow 1$ such that $\kappa$ measures the unit area. The susceptibility will then tend to zero and therefore the derivative of the susceptibility will increase within the trivial phase reaching zero when $M/B\rightarrow+\infty$.

### v2

@@ Abstract
Here, we elaborate on and develop the geometrical approach introduced in K. Le Hur, Physics Reports 1104 1-42 (2025) between the magnetic monopole created from a radial field, quantum physics and topological lattice models through quantum phase transitions. We introduce an effective magnetic moment for a monopole when applying an additional source field along z-direction which also mediates the quantum phase transition. We present its relation with the transverse pumped quantum Hall current. The magnetic susceptibility can be introduced as a measure of the topological invariant i.e. it remains quantized within the topological phase until the transition. We show the relation with two-dimensional topological lattice models such as a honeycomb Haldane model in real space. We develop the theory and present a numerical analysis between local invariants in momentum space introduced from Dirac points, correlation functions and the responses to circularly polarized light. We develop the formalism for coupled-planes materials including the possibility of quantum spin Hall effect and address a relation between the Ramanujan infinite alternating series and an interface in real space with a topological number one-half.
@@ 1 Introduction
The theory of magnetic monopoles attracts attention since the works of P. Curie [1] and P. A. M. Dirac [2, 3]. Detecting such monopoles associated to the planetary system, the universe, is yet of great interest. Recently, one of us has introduced a geometrical approach showing a correspondence between classical magnetic monopoles and quantum magnetic monopoles from a formulation at the poles on the sphere [4]. The magnetic monopole is produced from a radial magnetic field. Topological properties or equivalently the magnetic topological charge are generally measured through the integral of the magnetic field or through the integral of the Berry curvature in quantum physics on the surface area [5]. This approach [4] can be precisely rephrased into an effective cartesian metric in the vicinity of the poles in terms of a classical vector potential or in terms of the Berry gauge potential [5] in quantum physics associated to a coherent gauge (phase) representation for the eigenstates of the spin-1/2 particle on the Bloch sphere. Equivalently, the information on the magnetic charge is transported on each side of the equator to the poles in a thin handle. This formalism finds several practical applications [4]. It allows for a description of the quantum Hall response when building an analogy to the Newtonian approach. When applying an electric field along the polar angle an electron goes down from north to south pole with the induction of a transverse pumped current or a topological transverse pumped charge which is related to the quantum Hall conductivity and to the physics at the edges on a cylinder. It also allows us to show how photo-electric effects can become topological [6]. The responses to circularly polarized light can measure the topological charge or the quantum Hall conductivity from the poles which can be directly applied in two-dimensional (2D) topological lattice models through the Haldane model [7] and the quantum anomalous Hall effect (QAH) [8, 9] associated to the $\mathbb{Z}$ quantized invariant, where the poles correspond to the two Dirac points of the honeycomb lattice. From the same light signals, the quantum Hall conductivity [10] was first introduced through a summation on the wavevectors of the whole Brillouin zone [11, 12, 13]. Our geometrical formulation can also account for interaction effects in the detection of light in a quantitative manner through a variational stochastic approach [14, 15]. When adding interaction effects on the lattice in the Haldane model, our variational stochastic approach is very efficient to reveal the first-order nature of the quantum phase transition towards the Mott phase [14]. This approach is also adapted for two-dimensional topological insulators, i.e. the quantum spin Hall (QSH) effect and the Kane-Mele model [16, 17, 18], where circularly polarized light can also detect the $\mathbb{Z}_{2}$ invariant locally from the Dirac points[6]. It is important to emphasize that through the effect of a radial magnetic field magnetic monopoles are realized in quantum circuits with applications to the Haldane model [19, 20, 21]. The quantum metric tensor is also measured associated to phase transitions [21]. Within our approach, we have shown that circularly polarized light produces the same effect as a boost of the azimuthal angle building a correspondence between quantum metric and responses to circularly polarized light through the introduction of a geometrical function [4]. It is also possible to measure the geometrical information through the optical conductivity in quantum materials [22].
@@
What are the goals in this article ? We develop our geometrical approach towards quantum phase transitions, addressing transport properties and responses to circularly polarized light. The quantum phase transition is induced through a fixed magnetic field $M$ polarized along $z$ direction. In previous works, we have shown how the topological invariant can also be re-written in terms of the spin responses at the poles related to a protected quantum dynamo effect [23, 4, 24], with applications on topological quantum wires [25, 26, 27]. In this work, we evaluate the spin response on the surface of the ball, which is similar to a Skyrmion for $M=0$ as shown in Fig. 1. We introduce the effective magnetic moment of the monopole as a function of $M$ associated to the spin response in Eq. (20). We show how this is related to the transport properties. We introduce the associated susceptibility for the monopole which remains invariant within the topological phase until the phase transition and therefore can be introduced as a topological marker. To the best of our knowledge, this magnetic susceptibility of the monopole was not mentioned before in the literature. The magnetic moment has an orbital interpretation in the sense of the angular momentum associated to the transverse pumped current (charge) and this topological susceptibility shows then an analogy with orbital magnetism where the quantum Hall response is also thought of as a topological orbital susceptibility [28] corresponding to the derivative of the orbital magnetic moment with respect to the applied perpendicular magnetic field. We will then develop the theory for the Haldane model in two dimensions [7] with the introduction of a Semenoff mass [29] which corresponds e.g. to the effect of a charge density wave substrate [30]. Topological properties describing the occurrence of an edge mode are generally robust until the phase transition. Above the transition, particles progressively reside on one site preferably as a result of the proximity effect referring to a trivial insulator which does not present an edge mode and is characterized through a zero index (invariant). Finding observables to describe such a transition towards this trivial phase is yet of great interest related to the relative populations on the two sublattices. We develop an analysis of observables in momentum and real space. We take time to present a numerical analysis showing how our geometrical approach from the poles and from the Dirac points is particularly efficient. We emphasize that Berry curvatures at the two Dirac points encode information on the quantum Hall conductivity locally in momentum space [6]. It is relevant to mention here recent progress on measuring locally topological or geometrical properties in momentum space from photo-luminescence related to the Bloch sphere approach [31].
@@
The organization of our thoughts is as follows. In Sec. 2, we develop the geometrical aspects of the monopole in quantum physics associated to the quantum phase transition and we present the evaluation of the spin response and of the effective magnetic moment as a function of the magnetic field $M$. We also present the relation with the (mean) quantized transverse Hall current from a Newtonian approach. In Sec. 3, we develop the theory for the honeycomb Haldane model in the presence of a Semenoff mass producing the quantum phase transition. We present a precise map from the rhomboid Brillouin zone onto the sphere and derive analytically and numerically the local responses associated to local invariants in momentum (reciprocal) space. The spin-1/2 on the Bloch sphere will correspond to the pseudo spin-1/2 measuring the population imbalance between the two sublattices of the honeycomb lattice. We introduce a real space analysis related to the effective magnetic moment of the monopole. We elaborate on the local responses from circularly polarized light associated to the quantum phase transition. In Sec. 4, we develop the theory of coupled-planes related to the $\frac{1}{2}$ topological invariant building an analogy between a topological interface and the Ramanujan infinite alternating series. We analyse transport properties and the responses to circularly light within this correspondence. The coupled-planes model includes the alternating presence of QAH and QSH effects. In Sec. 5, we summarize our findings. Appendices are devoted to additional mathematical proofs. In Appendix A, we elaborate on the relation between the geometrical approach and transport, quantum Hall response with a map onto a cylinder. In Appendix B, we present a geometrical justification of the $\frac{1}{2}$ topological invariant for the coupled planes when reaching the infinity limit.
@@ 2 Bloch Sphere as a Magnetic Monopole, Quantum Hall Response and Topological Magnetism
In this Section, we study the physical responses as a function of the parameter $M$ corresponding to the fixed additional magnetic field along $z$ direction which will drive the quantum phase transition associated to the Poincaré-Hopf theorem. We evaluate the local spin magnetization responses at and in the vicinity of the poles on the Bloch sphere as a function of $M$. We present a simple proof in a dressed polar angle representation showing how the topological number is generally measurable locally through the spin response [23, 4, 24]. We will then introduce the mean (medium) spin magnetization on the surface area and address the relation with the transverse quantum Hall current or transverse pumped charge in a Newtonian (gravitational) correspondence when driving a charge from north to south pole through a longitudinal electric field. This will give rise to the effective topological magnetic moment $\kappa(M)$ in Eq. (20), associated to the loops of transverse currents formed along the equatorial plane on the whole surface. When $M=0$, as a result of the radial magnetic field, magnetic moments of the two hemispheres compensate each other and $\kappa(M=0)=0$. This analysis will also be related to the cylinder geometry [4].
@@
The last equation is equivalent to $|\alpha|^{2}=\cos^{2}\frac{\tilde{\theta}}{2}$ and $|\beta|^{2}=\sin^{2}\frac{\tilde{\theta}}{2}$. The topological invariant is measured for any fixed value of $M$. This local formulation of the topological invariant from the poles is gauge invariant and it agrees with general geometrical thoughts [4, 24] (see Appendix A). All the dependence on the parameter $M$ is then hidden into the definitions of the angles $\tilde{\theta}(\theta=0)$ and $\tilde{\theta}(\theta=\pi)$. We emphasize that the eigenstates are introduced within the same coherent gauge on the whole surface.
Figure 2: (Left) Classical magnetic monopole. We generalize the classical analogue of $A_{\varphi}$ introduced in Eq. (10) of Ref. [4] as $A_{\varphi}=-\frac{\cos\theta}{2}-\frac{M}{4}\cos(2\theta)$ (for a unit sphere). The component induced from the $M$ term in Eq. (10) does not modify the topological invariant. (Right) Representation of the effective topological properties for the quantum situation when including the magnetic field $M$ through the (dressed) angle $\tilde{\theta}(\theta=\pi)$. For all the phases the north pole is identified with $\theta=0$ i.e. $\tilde{\theta}=0$. Within the topological phase $B<M$, the south pole corresponds to $\tilde{\theta}(\theta=\pi)=\pi$. For $M>B$, the north and south poles are identical and the equivalent geometry does not encircle the topological charge at the origin. For $M=B$, geometrical properties are equivalent to a half sphere corresponding to a half monopole or a half Skyrmion.
As shown in Fig. 2, for all values of $M$, the north pole associated to $\theta=0$ is also equivalent to $\tilde{\theta}=0$ i.e. $\tilde{\theta}(\theta=0)=0$. On the other hand, we do observe a quantum phase transition when $M=B$ from the angle $\tilde{\theta}$ : as long as $M<B$, then $\tilde{\theta}=\pi$ also corresponds to south pole i.e. to $\theta=\pi$ such that $\tilde{\theta}(\theta=\pi)=\pi$, when $M=B$ the topological properties become equivalent to a half of a sphere $\tilde{\theta}(\theta=\pi)=\frac{\pi}{2}$ and for $M>B$ we also have the important equality $\tilde{\theta}(\theta=\pi)=\tilde{\theta}(\theta=0)$ implying $C=0$. The topological number then jumps from one when $M<B$ to zero when $M>B$. Within this formulation, at the quantum phase transition we also identify
@@
which can be interpreted as a half Skyrmion. It is interesting to mention that $A_{\varphi}(\tilde{\theta}(\theta=0))=+\frac{1}{2}$ corresponds to a topological characterization of a $\pi$ Berry phase encircling the north pole [24, 33].
From Ehrenfest’s theorem $\langle\psi|\sigma_{z}|\psi\rangle=\langle\sigma_{z}\rangle=\cos\tilde{\theta}$. Then, the topological number can be reformulated through the spin polarizations at the north pole and at the dressed angle associated to $\theta=\pi$ :
@@
This definition agrees with previous results on one sphere when $M=0$ [23]. This formula was adapted for the analysis of topological quantum wires with applications in real space associated to correlation functions [25, 26, 27]. The profile of the magnetic structure at the poles in Fig. 1 clearly reveals the quantum phase transition. At the topological quantum phase transition, we have the identity
@@
For two spheres, it is possible to reach fractional $\frac{1}{2}$ numbers on a line or in a phase of the parameters space associated to the interaction between spins [4]. For $M>B$, the spin magnetizations associated to $\theta=0$ and $\theta=\pi$ are the same such that $C=0$ i.e. $\tilde{\theta}(\theta=0)=\tilde{\theta}(\theta=\pi)$.
@@
The charge $e$ is the charge of an electron. The pumped charge at an angle $\tilde{\theta}$ reads
@@
In this formulation, in the presence of the radial magnetic field acting on the spin of an electron and in the presence of the magnetic field source $M$, we measure the transverse pumped charge associated to an electric field applied along the polar angle such that $\hbar\dot{\tilde{\theta}}=eE$. At time $t=0$, a charge is present at north pole. The charge goes down similar to the apple and the gravitational force is a result of the electric field directed along the polar angle direction i.e. corresponds to a Coulomb force. The electron moves on the surface of this metallic ball associated to the Bloch sphere. The final time $T$ is fixed to be identical for any $M$ such that the angle $\tilde{\theta}$ reaches $\pi$ i.e. $T=\frac{eE}{\hbar\pi}=\frac{2eE}{h}$. Within the topological phase, the transverse pumped charge at south pole precisely measures the topological invariant itself $C=1$. When we reach the transition maintaining the same value of the (final) time $T$ for the measure, associated to Fig. 2, this corresponds to maintain the Berry gauge potential fixed to $A_{\varphi}(\tilde{\theta}=\pi)=A_{\varphi}(\tilde{\theta}=\frac{\pi}{2})=0$ for dressed angles $\tilde{\theta}\in[\frac{\pi}{2},\pi]$ such that there is no Berry curvature effectively present in the south hemisphere. In that case, the topological invariant becomes halved i.e. equal to $C_{1/2}$ and the transverse pumped charge is also halved. Above the transition, the charge remains at the north pole i.e. $A_{\varphi}(\tilde{\theta}=\pi)=A_{\varphi}(\tilde{\theta}=0)$ and there is no transverse pumped current. In Appendix A, we also show the relation with the cylinder geometry.
@@
Here, $\kappa$ is the mean of the spin response or the effective spin response as a function of $M$ associated to the surface area in Fig. 1. When $M=0$, from symmetry between the two hemispheres $\kappa=0$. In response to the radial magnetic field in Fig. 1 the two hemispheres produce an effective moment equal in amplitude but opposite in directions. We are precisely questioning the (topological) properties of $\kappa$ as a function of $M$ when inserting the equation
@@
When $M\rightarrow+\infty$ this results in $\langle\sigma_{z}\rangle\rightarrow 1$ such that $\kappa$ measures the unit area. The susceptibility will then tend to zero.